OpenStudy (anonymous):

Identify which of the following functions decreases as x increases.: f(x)=log_4/5^x or g(x)=log_3^x

hero (hero):

Can you write that out? It is not clear what the bases are.

OpenStudy (anonymous):

okay \[f(x)=\log _{4/5^{X}}\] or \[g(x)= \log _{3^{X}}\]

OpenStudy (anonymous):

not sure how to do the fraction thing on the equation function on here but the 4/5 is a fraction on the first one.

hero (hero):

Hint: The one of the bases is a whole number; the other base is a fraction.

OpenStudy (anonymous):

so the answer is the second equation because it's not a fraction? lol just guessing. i think i had a question like this before.

hero (hero):

Actually, they want the function that DECREASES as x increases

hero (hero):

logs of fractions produce negatives

OpenStudy (anonymous):

oh okay. so since negatives are going lower, the answer would be the first one?

hero (hero):

:D

OpenStudy (anonymous):

thanks honey <3